10. Oktober 2003. U. Schoenwaelder; http://www.math.rwth-aachen.de/∼Ulrich.Schoenwaelder HB = Hochschulbibl. RWTH, HBZ = http://www.hbz-nrw.de/ (HBZ-CD-ROM Online), MB = Mathematikbibl., DB = Didaktikbibl. (Winter), EB = Erziehungswiss. Bibl., FH = Bibl. Fachhochschule Aachen, FL = Fernleihe, IB = Informatikbibliothek, IB Nr. = Institutsbibliothek Nr., LB = HB– Lehrbuchsammlung, LS = HB–Lesesaal

Literatur zur Algebra I

[1] Sukumar Das Adhikari. An Introduction to Commutative Algebra and Number Theory. New Delhi: Narosa; Chapman & Hall/CRC, 1999, 2002. ISBN 81-7319304-5; 0-8493-0990-5. Elementary introduction. Groups, rings, fields, polynomial rings, Euclidean domains, field extensions, modules.

[2] W. A. Adkins and S. H. Weintraub. Algebra. An Approach via Module Theory. GTM 136. Springer–Verlag, 1992. MB: 16528, HB: Bb 1244-136.

[3] R. B. Allenby. Rings, Fields and Groups: an Introduction to Abstract Algebra. Arnold, London, ^{1}1983, ^{2}1991. MB: 11912, 15989.

[4] M. Anderson and T. Feil. A First Course in Abstract Algebra: Rings, Groups, and Fields. PWS, 1995. Tel. Rev.: Monthly 102:8 (95), 759: One of a few nice texts that begin with rings.

[5] M. Artin. Algebra. Birkhäuser Advanced Texts. Basler Lehrbücher. Birkhäuser, 1993, ^{2}1998. ISBN 3-7643-5938-2. MB: 16135, HB: Bb 1919. Auch Symmetrie, lineare Gruppen, quadratische Zahlkörper.

[6] B. Artmann. Eine Einführung in die Algebra. Unter ständiger Berücksichtigung der ganzen und der rationalen Zahlen. Moderne Mathematik in elementarer Darstellung 13. Vandenhoeck & Ruprecht, 1973. MB: 8561, HB: 13 Bb 1126.

[7] J. A. Beachy and W. D. Blair. Abstract Algebra. Waveland Press, Prospect Heights, Illinois, 60070, ^{2}1996. MB: Info at http://www.math.niu.edu/˜ beachy/abstract algebra.

[8] G. Böhme. Algebra. Anwendungsorientierte Mathematik. Springer–Lehrbuch. Springer–Verlag, ^{7}1992. MB: 16436, HB: Bb 1118-1+3, auch Lehrbuchsammlung.

[9] S. Bosch. Algebra. Springer–Lehrbuch. Springer–Verlag, 1993, ^{2}1996. MB: 17225, HB: Bb 1903. Auch Resultanten, Diskriminanten und symmetrische Funktionen, Tensorprodukt.

[10] N. Bourbaki. Algebra I (Chapters 1–3). Originally published as Éléments de Mathématique, Algèbre 1–3 (1970). Springer– Verlag, 1974, ^{2}1988. MB: 8636a, HB: 2,1 Bb 1301. ISBN 3-540-64243-9.

[11] Rolf Brandl. Alle Algebra-Aufgaben ... die Lösungen. Verlag Rolf Brandl, Postfach 3045, D-95006 Hof, 2000. HBZ. Vgl. [12].

[12] Rolf Brandl. Alle Algebra-Aufgaben: ein Repetitorium über Gruppen, Ringe, Körper und Galois-Theorie. Verlag Rolf Brandl, Postfach 3045, D-95006 Hof, 2000. HBZ. Vgl. [11].

[13] Oscar Campoli. A principal ideal domain that is not a euclidean domain. Amer. Math. Monthly, 95:868–871, 1988. MB: Z 42. The author shows that Z[θ] for θ = (1 + \sqrt{−19})/2 is a PID that is not a Euclidean domain.

[14] L. Childs. A Concrete Introduction to Higher Algebra. UTM. Springer–Verlag, ^{1}1979,^{5}1992. MB: 10252. HB: Bf 7083+2.

[15] Arjeh M. Cohen, Hans Cuypers, and Hans Sterk. Algebra Interactive! Learning Algebra in an Exciting Way. Springer– Verlag, 1999. ISBN 3-540-65368-6. HB: BB5081.

[16] P. M. Cohn. Algebra (Volumes 1 – 3). John Wiley & Sons, ^{2}1995. HB: Bf 6126-3+2.

[17] P. M. Cohn. Algebra (Volumes 1 – 2). John Wiley & Sons, Vol. 1: ^{2}1982, Vol. 2: ^{2}1989. MB: 12191a, b. Vol. 1: §4.8: Affine spaces. Vol.2: §1.3: Categories, §4.5: Projective and injective modules, §4.7: The tensor product of modules, §8.3: The p–adic numbers, §9.5: Dedekind domains. Lebensdaten im Index.

[18] R. Dedekind. Theory of Algebraic Integers. Cambridge Mathematical Library. Cambridge University Press, 1996. MB: 17 991. Part I: Translators Introduction (pp. 1–50); Part II: Translation by John Stillwell of: R. Dedekind, Sur la Théorie des Nombres Entiers Algébriques, Gauthiers–Villars, 1877 (pp. 51–152).

[19] E. Dubinsky and U. Leron. Learning Abstract Algebra with ISETL. With 1 IBM-PC floppy disk (3.5 inch). Springer– Verlag, 1994. HB: BN1921.

[20] D. S. Dummit and R. M. Foote. Abstract Algebra. Prentice Hall, 1991, ^{2}1999. ISBN 0-471-36857-1. Zbl. 751.00001, pre991.27642. Tel. Rev.: Amer. Math. Monthly 100(1):1993, 96: Ample content and sophistication for a year course for advanced undergraduates.

[21] John R. Durbin. Modern Algebra: an Introduction. Wiley, Chichester, England, ^{3}1992. ISBN 0-471-51001-7.

[22] Graham Ellis. Rings and Fields. Oxford Univ. Press, 1992. ISBN 0-19-853455-8, 0-19-853454-X. HBZ.

[23] M. H. Fenrick. Introduction to the Galois Correspondence. Birkhäuser, 1991. MB: 16428.

[24] Benjamin Fine and Gerhard Rosenberger. The Fundamental Theorem of Algebra. UTM. Springer–Verlag, 1997. ISBN 0-387-94657-8. HBZ.

[25] G. Fischer and R. Sacher. Einführung in die Algebra. Teubner, ^{3}1983. MB: 9635, HB: Bf6194. Enthält einen Großteil des Vorlesungsstoffes.

[26] John B. Fraleigh. A First Course in Abstract Algebra. Addison-Wesley series in mathematics. Reading, MA: Addison-Wesley, ^{2}1976. ISBN 0-201-01984-1. HB: Bb6591+2.

[27] J. A. Gallian. Contemporary Abstract Algebra. D. C. Heath and Company, ^{3}1994. MB: 17219. Lst 1656. Ch. 28: Frieze Groups and Crystallographic Groups (ohne Definition der geometrischen Äquivalenz). Vita vieler Mathematiker. Suggested readings.

[28] I. M. Gelfand and A. Shen. Algebra. Birkhäuser, Boston, 1993. ISBN 0-8176-3677-3, 3-7643-3677-3. HBZ.

[29] L. Gerritzen. Grundbegriffe der Algebra. Eine Einführung unter Berücksichtigung funktorieller Aspekte. Vieweg, 1994. HB: Bb 1927.

[30] J. K. Goldhaber and G. Ehrlich. Algebra. Macmillan, London, 1970. MB: 7109.

[31] Pierre Antoine Grillet. Algebra. Pure and Applied Mathematics. Wiley, 1999. MB: 18762. ISBN 0-471-25243-3. Umfangreich. Groups; Rings; Fields; Rings and Modules; Other Structures (Categories, Universal Algebra, Lattices); Appendices (Numbers, Sets and Order.

[32] I. N. Herstein. Algebra. Übersetzung von Topics in Algebra, ^{2}1977. Physik–Verlag, 1978. MB: 9911.

[33] I. N. Herstein. Abstract Algebra. Macmillan, 1986, ^{2}1990. MB: 15662. Enthält das Allerwichtigste.

[34] I. N. Herstein. Topics in Algebra. Wiley, Chichester, England, ^{2}1977. MB: 6211 (1964).

[35] B. Hornfeck. Algebra. Walter de Gruyter & Co., 1969. MB: 4588, HB: Bf5251.

[36] I. M. Isaacs. Roots of polynomials in algebraic extension fields. Amer. Math. Monthly, 87(7):543–544, 1980. MB: Z 42.

[37] I. M. Isaacs. Algebra: a Graduate Course. Brooks/Cole Publ. Co., 1994. ISBN 0-534-19002-2. MB: 19683. HBZ. MR 95k: 00003.

[38] N. Jacobson. Lectures in Abstract Algebra. I – Basic Concepts. D. van Nostrand, 1951, ... MB: 913 c, HB: Bf54-1+1.

[39] N. Jacobson. Basic Algebra I, II. W. H. Freeman, 1980. MB: 11603 a, b.

[40] A. Jones, S. A. Morris, and K. R. Pearson. Abstract Algebra and Famous Impossibilities. Universitext. Springer–Verlag, 1991. MB: 16328.

[41] H. Kaiser, R. Lidl, and J. Wiesenbauer. Aufgabensammlung zur Algebra. Studienbuch für Studierende der Mathematik und Physik ab 1. Semester. studien–texte Mathematik. Akademische Verlagsgesellschaft, Wiesbaden, 1975. MB 10039.

[42] George R. Kempf. Algebraic Structures. Vieweg, 1995. ISBN 3-528-06583-4. HBZ viele.

[43] F. Klein. Vorlesungen über das Ikosaeder und die Auflösung der Gleichungen vom fünften Grade. Herausgegeben mit einer Einführung und Kommentaren von P. Slodowy. Birkhäuser, 1992. ISBN 3-8154-2032-6. MB: 4165. HBZ.

[44] Dennis Kletzing. Abstract Algebra. Harcourt Brace, 1991. ISBN 0-15-500391-7. HBZ. Algebraic concepts are introduced geometrically.

[45] Klin, Pöschl, and Rosenberg. Angewandte Algebra für Mathematiker und Informatiker: Einführung in gruppentheoretischkombinatorische Methoden. Berlin: Dt.Verl. d. Wiss.; Braunschweig: Vieweg, 1988. ISBN 3-326-00190-8; 3-528-08985-7. MB: 14333. HBZ. Abzähltheorie.

[46] O. Koerner. Algebra. Studienbuch für Studierende der Mathematik und Physik ab 2. Semester. Studien–Text. Frankfurt am Main: Akad. Verlagsges.; Wiesbaden: Aula-Verl., 1974, ^{2}1990. ISBN 3-400-00269-0; 3-89104-504-2. HB: Bf6158, MB: 7841 1974; HBZ.

[47] Ernst Kunz. Algebra. Vieweg–Studium: Aufbaukurs Mathematik 43. Vieweg, Braunschweig, 1991. ISBN 3-528-07243-1. MB: 16021, HB: Bf9311, Bf9311+2 (1994). HBZ. Wohlmotivierte Einführung in Körper– und Galois–Theorie.

[48] T. Y. Lam. A First Course in Noncommutative Rings. GTM 131. Springer–Verlag, 1991. MB: 16008. HB: Bb 1244-131. Ch. 5: Introduction to Division Rings.

[49] T.-Y. Lam. Lectures on Modules and Rings. GTM 189. Springer, 1999. ISBN 0-387-98428-3. HBZ.

[50] S. Lang. Undergraduate Algebra. UTM. Springer–Verlag, 1987,^{2}1990. ISBN 0-387-97279-X. MB: 13605. HBZ.

[51] S. Lang. Algebra. World Student Series. Addison–Wesley, 1965, ^{2}1984, ^{3}1993. MB: 12409, 16801, HB: Bf5068.

[52] S. Lang. Algebraic Structures. Addison-Wesley Series in Mathematics. Addison-Wesley, 1967, ^{2}1968. ISBN 0-201-04173-1. HB: BF5553. Die wichtigsten algebraischen Strukturen knapp und übersichtlich.

[53] S. Lang. Algebraische Strukturen. Moderne Mathematik in elementarer Darstellung. Vandenhoeck & Ruprcht, 1979. ISBN 3-525-40542-1. HB: BB1126-18+1. Die wichtigsten algebraischen Strukturen knapp und übersichtlich.

[54] S. Lang. Algebra. Third Edition. World Student Series. Addison–Wesley, 1993. ISBN 0-201-55540-9. MB: 16801, HB: Bf 5068+3, LB und LS. Umfassend.

[55] M. A. Lee. Some irreducible polynomials which are reducible mod p for all p. Amer. Math. Monthly, 76:1125, 1969. MB: Z 42. This brief note gives a class of polynomials that are reducible mod p for all primes p, but are irreducible over the integers.

[56] R. Lidl. Algebra für Naturwissenschaftler und Ingenieure. Sammlung Göschen 2120. Walter de Gruyter, Berlin etc., 1975. ISBN 3-11-004729-2. MB 8541, HB: Bf 6464. Grundlagen.

[57] R. Lidl and G. Pilz. Angewandte abstrakte Algebra I, II. B.I.–Wissenschaftsverlag, Mannheim etc., 1982. ISBN 3-411-01620-5. MB 11543 a, b. HB: BB1544-1+1, BB1544-2+1. Literaturangaben zu den einzelnen Kapiteln.

[58] R. Lidl and G. Pilz. Applied Abstract Algebra. Undergraduate Texts in Mathematics. Springer–Verlag, New York etc., 1984. ISBN 0-387-98290-6. MB 12609. HBZ. Boolesche Algebren und Verbände, endliche Körper und Polynome (Berlekamp– Algorithmus), Codes, RSA, Pólyas Abzähltheorie, Automaten, formale Sprachen.

[59] R. Lidl and J. Wiesenbauer. Ringtheorie und Anwendungen. Grundlagen und Anwendungsbeispiele in der Kodierungstheorie und in der Genetik. studien–texte Mathematik. Akademische Verlagsgesellschaft, Wiesbaden, 1980. ISBN 3-400-00371-9. MB: 11542. HBZ. Rechnen in/über endlichen Körpern. Berlekamp–Algorithmus S. 177.

[60] F. Lorenz. Einführung in die Algebra I, II. Mannheim: BI, ^{3}1992. MB: 15680 a,b, HB: Bf930-1+2, Bf930-2+1. Günstige LS–Konvention.

[61] H. Lüneburg. Einführung in die Algebra. Hochschultext. Springer–Verlag, Berlin etc., 1973. MB: 7522, HB: Bf 6004 LB und LS.

[62] S. MacLane and G. Birkhoff. Algebra. Macmillan, London, 1967, ^{2}1979. MB: 10555, HB: Bf 5248 und LB, LS.

[63] D. S. Malik, John M. Mordeson, and M. K. Sen. Fundamentals of Abstract Algebra. Intern. series in pure & applied mathematics. McGraw–Hill, 1997. ISBN 0-07-040035-0. HB: BB587. Contents: Sets, Relations and Integers / Intoduction to Groups / Permutation Groups / Subgroups and Normal Subgroups / Homomorphisms and Isomorphisms of Groups / Sylow Theorem / Solvable and Nilpotent Groups / Finitely Generated Abelian Groups / Introduction to Rings / Subrings / Ideals and Homomorphisms / Extensions of Rings / Direct Sum of Rings / Polynomial Rings / Euclidean Domains / Unique Factorisation Domains / Maximal, Prime and Primary Ideals / Noetherian and Artinian Rings / Modules and Vector Spaces / Matrix Rings / Field Extensions / Algebraic Extensions / Multiplicity of Roots / Finite Fields / Galois Theory and Applications / Normal Extensions / Geometric Constructions / Binary Codes / Groebner Bases.

[64] J. I. Manin. Über die Lösbarkeit von Konstruktionsaufgaben mit Zirkel und Lineal. In P. S. Alexandroff et al., editor, Enzyklopädie der Elementarmathematik, Band IV Geometrie, Hochschulbücher für Mathematik Band 10, pages 203–228. VEB Deutscher Verlag der Wissenschaften, Berlin, 1969. MB: 6215 a, HB: 4^{1} Bb 1054. Übers. a. d. Russ.

[65] M. Marcus. Introduction to Modern Algebra. Monographs and Textbooks in Pure and Applied Mathematics 47. Dekker, 1978. MB: 9969. Ch. 4: Modules and Linear Operators; 4.1: The Hermite Normal Form, 4.2: The Smith Normal Form, 4.3: The Structure of Linear Transformations.

[66] K. Meyberg. Algebra, Teil 1. Carl Hanser Verlag, München, Wien, 1975. MB: 8048 a. HB: Bf 6227-1+1 und LB.

[67] K. Meyberg. Algebra, Teil 2. Carl Hanser Verlag, München, Wien, 1976. MB: 8048 b. HB: Bf 6227-2+1 LB und LS.

[68] T. T. Moh. Algebra. Series on University Mathematics 5. World Scientific, 1992. ISBN 981-02-1196-1. HBZ.

[69] M. Nagata. Theory of Commutative Fields. Translations of Mathematical Monographs 125. AMS, 1993. ISBN 0-8218-4572-1. MB: 16874. HBZ.

[70] P. M. Neumann, G. A. Stoy, and E. C. Thompson. Groups and Geometry. Oxford Univ. Press, 1994. ISBN 0-19-853451-5.

[71] A. L. Oniščik and R. Sulanke. Algebra und Geometrie, Band 1: Eine Einführung. VEB DVW, 1986. MB: 13480 a, HB: Bf 7908-1+1.

[72] A. L. Oniščik and R. Sulanke. Algebra und Geometrie, Band 2: Moduln und Algebren. VEB DVW, 1988. MB: 13480 b, HB: Bf 7908-2+1.

[73] V. Pless. Cyclotomy and cyclic codes. In A. Burr et al., editor, The Unreasonable Effectiveness of Number Theory, Proceedings of Symposia in Applied Mathematics. AMS, 1993.

[74] C. Procesi. Elementi di teoria dei gruppi. Decibel Editrice, Padova; Nicola Zanichelli Editore S.p.A., Bologna, ^{3}1990. MR 92a: 20001.

[75] L. Rédei. Algebra. Mathematik und ihre Anwendungen in Physik und Technik, R. A. 26. Geest & Portig, Leipzig, 1959. HB: 26 Za 5810.

[76] L. Rédei. Algebra I. Pergamon, Oxford, 1967. MB: 1524.

[77] H.-J. Reiffen, G. Scheja, and U. Vetter. Algebra. BI-Hochschultaschenbücher 110 a, 110. Mannheim: BI, 1969, ^{2}1984. ISBN 3-411-00110-0. MB: 4929, HB: ZA5069-110.

[78] J. Rotman. Galois Theory. Universitext. Springer–Verlag, 1990. ISBN 3-540-97305-2, 0-387-97305-2 . MB: 15404, HB: Bb 1784.

[79] W. W. Sawyer. Eine konkrete Einführung in die abstrakte Algebra. BI-Hochschultaschenbücher 492/492a. BI, 1959. MB: 5818.

[80] John Scherk. Algebra: a Computational Introduction. Chapman & Hall/CRC, 2000. ISBN 1-58488-064-3. MB: 19457. Basic group and Galois theory.

[81] D. Scott and D. R. Peeples. A constructive proof of the partial fraction decomposition. American Math. Monthly, 95(7):651–653, 1988. MB: Z 42. Verkappter Chinesischer Restsatz?

[82] I. R. Shafarevich. Algebra I: Basic Notions of Algebra. Encyclopedia of Mathematical Sciences 11. Springer–Verlag, 1990. MB: 15143 a (Verlustliste 1996). HB: Bb 1690-11+1. Übersichtsartig.

[83] D. Sharpe. Rings and Factorization. Cambridge Univ. Press, 1987. HB: Bf 907. Euklidische und Gaußsche Bereiche und Anwendungen.

[84] C. C. Sims. Abstract Algebra. A Computational Approach. John Wiley & Sons, New York etc., 1984. MB: 12718.

[85] Sahib Singh. Non–euclidean domains: an example. Mathematics Magazine, 49:243, 1976. MB: Z 167. This note gives a short proof that Z[\sqrt{−n}] is an integral domain that is not euclidean when n > 2 and −n = 2 or 3 modulo 4.

[86] K. Spindler. Abstract Algebra with Applications. Volume I: Vector Spaces and Groups, Volume II: Rings and Fields. Dekker, 1994. ISBN I: 0-8247-9144-4, II: 0-8247-9159-2k. MB: 19235. In Vol. I: Tensor Products and Base Field Extension.

[87] I. Stewart. Galois Theory. Chapman and Hall, 1973, ^{2}1989, ^{3}2003. MB: 7107. HBZ. Historical Introduction and Overview added. 3rd edition: now begins with polynomials over the complex numbers, contains several newer proofs, four proofs for insolubility of the quintic.

[88] J. Stillwell. Mathematics and Its History. UTM. Springer–Verlag, 1989. MB: 14970. Ch. 13: Complex Numbers in Algebra (Cubic Equations). S. 126: Generating Functions, Zeta Function. Bibliographical Notes: Pythagoras (S. 11), Diophantus (S. 36), Archimedes (S. 46), Tartaglia, Cardano, Viète (S. 59), Descartes (S. 74), Desargues, Pascal (S. 96), Wallis, Newton, Leibniz (S. 110), Gregory, Euler (S. 130), Fermat (S. 149), Abel, Jacobi (S. 161), The Bernoullis (S. 181), Riemann (S. 215), Lagrange, Cauchy (S. 232), Harriot, Gauss (S. 250), Bolyai, Lobachevsky (S. 271), Galois (S. 287), Poincarè (S. 309), Gödel (S. 328).

[89] J. Stillwell. Elements of Algebra; Geometry, Numbers, Equations. Undergraduate Texts in Mathematics. Springer–Verlag, 1994, 1996. ISBN 0-387-94290-4. HBZ. Developing algebra out of classical number theory and geometry.

[90] B. L. van der Waerden. Algebra I. Heidelberger Taschenbücher 12. Springer–Verlag, Berlin, ^{4}1955, ^{8}1971. MB: 518, HB: Bf 1183-1+8. Weitere Auflagen: ^{9}1993.

[91] B. L. van der Waerden. Algebra. Zweiter Teil. Heidelberger Taschenbücher 23. Springer–Verlag, Berlin etc., ^{5}1967. MB: 3834, HB: Bf 1183-2+5 und LB. Weitere Auflagen: ^{6}1993.

[92] E. Walker. Introduction to Abstract Algebra. The Random House/Birkhäuser Series. Random House, 1987. FL: UB Paderborn P41 TDP 3313. Abstrakte Theorie betont.

[93] Gisbert Wüstholz. Grundlagen der Algebra. Für Studierende der Mathematik, Physik, Informatik. Vieweg, 2004. ISBN 3-528-07291-1. Ring- und Körpertheorie, Darstellungstheorie von endlichen Gruppen, Algebren, Moduln.